OpenStudy (anonymous):

An exponential function is written as F(x) = a x b^x, where the coefficient a is a constant, the base b is _____ but not equal to 1, and the exponent x is any number. A. an integer B. negative C. real D. positive

OpenStudy (triciaal):

usually positive whole number

OpenStudy (anonymous):

So D?

OpenStudy (triciaal):

never thought about it sounds good to me

OpenStudy (triciaal):

@xapproachesinfinity , @jim_thompson5910 can we have a negative base for exponential functions?

jimthompson5910 (jim_thompson5910):

you can, but only if the exponent is an integer. You run into problems if the exponent is a fraction or decimal number

jimthompson5910 (jim_thompson5910):

so it's usually a good idea to make the base positive

OpenStudy (anonymous):

so stay with D?

jimthompson5910 (jim_thompson5910):

yeah stay positive

OpenStudy (xapproachesinfinity):

i will always see a^x where a>0

OpenStudy (triciaal):

thanks guys

OpenStudy (freckles):

y=b^x has to have b as positive consider b=-1 y=(-1)^x will not exist for all real x because example: (-1)^(1/2) is not real y=(-1)^x does not fit the definition of exponential function

OpenStudy (xapproachesinfinity):

yeah a<0 would complicate things just taking extreme cases would prove that

OpenStudy (anonymous):

Sweet thanks :D

OpenStudy (mertsj):

Definition of Exponential Function An Exponential Function is a function of the form y = ab^x, where both a and b are greater than 0 and b is not equal to 1.

jimthompson5910 (jim_thompson5910):

'a' can be negative though for y = a*b^x

OpenStudy (triciaal):

@jim_thompson5910 " where both a and b are greater than 0 " so 'a' canNOT be negative

jimthompson5910 (jim_thompson5910):

with y = a*b^x, only b has to be positive. The value of 'a' can be any real number you want.